Find the area under the graph of the function over the interval given 1 y= (1.4) 1.4 OA. 2 1 O B. OC. 4 OD 1
Answer #1

Given y = 4 / 4 y = -1 Area of the shaded region is given by; (415) A= s da (90) x=1 4 U=4 J 2/2 da - [2 x2] = 2(54-5) = 2(2-